# Is Contrastive Distillation Enough for Learning Comprehensive 3D Representations?

### v1

@@ Abstract
journal:
@@ 1 Introduction
LiDAR sensors have become essential tools for capturing detailed 3D information of the environment, playing a critical role in applications such as autonomous driving, robotics, and urban planning (Zhang et al., 2023b, 2024b, a). The rich geometric data from LiDAR point clouds provide valuable spatial awareness that is difficult to achieve with traditional 2D sensors. However, processing these 3D point clouds often requires vast amounts of labeled data to train deep neural networks effectively. Annotating point cloud data, however, is both time-consuming and expensive, posing significant challenges to the scalability and practicality of 3D deep learning models (Sautier et al., 2022; Chen et al., 2024). To address this issue, self-supervised learning has emerged as a promising solution, where networks are first trained on large volumes of unlabeled data (He et al., 2022; Caron et al., 2021). This pre-training process enables the model to learn useful feature representations without the need for costly manual annotations. Once pre-trained, the network can be fine-tuned on smaller labeled datasets, greatly reducing the need for extensive labeling efforts and improving the efficiency of training (Chen et al., 2020).
@@ 2 Related Work
3D Scene Understanding. Traditional methods for 3D scene understanding often rely on representations such as raw points (Choe et al., 2022; Chen et al., 2022a), voxels (Puy et al., 2023; Zhang et al., 2024d), range views (Kong et al., 2023a; Tian et al., 2022), and multi-view fusion (Fadadu et al., 2022; Xu et al., 2021) to capture environmental features. While effective, these approaches typically depend on large volumes of labeled data, which is both time-consuming and expensive to obtain, thereby hindering the scalability of 3D perception models (Liu et al., 2022b). To mitigate this issue, recent research has explored alternatives to reduce the reliance on fully annotated datasets. These include semi-supervised (Kong et al., 2023b; Ho et al., 2024), weakly-supervised (Liu et al., 2023; Chibane et al., 2022), self-supervised, and active learning (Luo et al., 2023; Xie et al., 2023) strategies, all of which aim to leverage minimal labeled data or automatically generate useful annotations to enhance model performance and scalability.
3D Representation Learning. Recent advancements in self-supervised learning for 3D point clouds have evolved alongside improvements in image-based methods (Zhang et al., 2024a, c). These include pretext tasks like predicting transformations or reconstructing point cloud parts (Poursaeed et al., 2020; Sauder and Sievers, 2019). Discriminative methods focus on contrastive learning across different levels of representation (point, segment, region) to capture geometric and structural information (Nunes et al., 2022; Yin et al., 2022; Chen et al., 2022b). Temporal-consistency methods leverage spatiotemporal correlations, aligning objects over time for robust representations (Nunes et al., 2023; Huang et al., 2021). Reconstruction-based methods, such as those using Chamfer distance or surface reconstruction, aim to recover point cloud details from masked data. Cross-modal distillation approaches utilize synchronized camera-LiDAR data for contrastive learning, transferring knowledge from 2D to 3D networks (Sautier et al., 2022; Mahmoud et al., 2023; Liu et al., 2024; Chen et al., 2024; Liao et al., 2024; Puy et al., 2024; Xu et al., 2025). In this paper, we identify the limitations of these cross-distillation methods theoretically and propose a novel framework for more comprehensive 3D representation learning.
@@ 3.1 Preliminary
[…]
@@ 4.1 Overview
Notation. Let $P=\{p_{1},p_{2},...,p_{N}|p_{i}\in\mathbb{R}^{3}\}$ be a point cloud consisting of $N$ points collected by a LiDAR sensor, and =$\mathcal{I}=\{I_{c}\|\ c=1,...,N_{\mathrm{cam}}\}$ multi-view images captured by $N_{\mathrm{cam}}$ synchronized cameras, where $I_{c}\in\mathbb{R}^{H\times W\times 3}$ is a single image with height $H$ and width $W$.
@@ 4.2 Multi-modal Unified Codebook
To map the extracted modality-shared features into a unified latent space, we adopt a vector quantization (VQ) mechanism (Van Den Oord et al., 2017). This approach discretizes the continuous representations into a finite set of codewords, ensuring consistent semantic alignment across different modalities. Specifically, we define an embedding table $E=\{e_{1},e_{2},\dots,e_{V}\}\in\mathbb{R}^{V\times D_{c}}$, where $V$ is the size of the codebook, and $D_{c}$ is the dimensionality of each codeword $e_{v}$.
@@
$\mathcal{L}_{\text{commit}}=\frac{1}{n^{\mathrm{3D}}_{v}(t)}\left(\sum_{i=1}^{n^{\mathrm{3D}}_{v}(t)}||F^{\mathrm{2D}}_{i}-\mathrm{sg}[e_{v^{*}}]||_{2}^{2}+||F^{\mathrm{3D}}_{i}-\mathrm{sg}[e_{v^{*}}]||_{2}^{2}\right),$
where $v^{*}=\underset{k\in\{1,\dots,V\}}{\mathrm{argmin}}\,||F^{\mathrm{3D}}_{i}-e_{k}||_{2}$.
@@ 4.5 Overall Objective Function
where reconstruction loss $\mathcal{L}_{\text{rec}}$ is the loss for the MIM task, orthogonal loss $\mathcal{L}_{\text{orth}}$ is designed to encourage independence between the modality-shared and modality-specific features within each modality, and $\mathcal{L}_{\text{kl}}$ is the Kullback-Leibler (KL) Divergence loss developed for semantic consistency proposed in OLIVINE (Zhang and Hou, 2024).
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm, which computes the squared sum of all entries in the matrix.
@@ 5.1 Implementation Details
Pre-training Details. We pre-train the model for 50 epochs with an initial learning rate of 0.001, which is adjusted using a one-cycle learning rate scheduler (Smith, 2017) and the Adam optimizer. The 2D backbone is a Vision Transformer (ViT) pretrained with MoCoV2 (Chen et al., 2020). The pre-training tasks are performed on four RTX 3090 GPUs, with a batch size of sixteen. During pretraining, each image input is randomly masked, with 50% of its content masked at each step. For occupancy estimation, we randomly select two thousand query points per scene. For fine-tuning, we adhere to the same data splits, augmentation strategies, and evaluation protocols as those used in prior works (Mahmoud et al., 2023; Chen et al., 2024) on the nuScenes and SemanticKITTI datasets, and apply a similar procedure on other datasets.
@@ 5.5 Ablation Study
Table 6: Ablation study of each component pre-trained and fine-tuned on nuScenes. PP: Basic pixel-point contrastive learning. Rec.: Image reconstruction and occupancy estimation tasks. Codebook: The multi-modal unified codebook. Geo.: Geometry-enhanced masked image modeling.
Exp. PP Rec. Codebook Geo. $\mathcal{L}_{\text{kl}}$ nuScene S.K.
@@ B. Implementation Details
Network Architectures. In our pre-training pipeline, the input images are resized to 416$\times$224. For the 3D semantic segmentation task, we utilize the Sparse Residual 3D U-Net 34 (SR-UNet34) (Ronneberger et al., 2015), following the methodology outlined in SLidR (Sautier et al., 2022). The SR-UNet34 outputs a feature map with 256 channels, while the image branch produces a 64-dimensional feature vector. To match the dimensionality of these features, we employ a 3D convolutional layer in the projection head to reduce the point feature map to 64 channels. For input, the 3D point data is converted into voxels, with Cartesian coordinates spanning an X-Y range of [-51.2m, 51.2m] and a Z-range of [-5.0m, 3.0m]. Each voxel has dimensions of (0.1m, 0.1m, 0.1m). For 3D object detection, we adopt VoxelNet (Zhou and Tuzel, 2018), with a maximum of 10 points per voxel and a limit of 60,000 voxels to process the point cloud input.
@@ C. Additional Experimental Results
Additional Quantitative Results. Tables 8 and 9 present the per-class performance of various point cloud pretraining methods, including our approach and several baseline models, all fine-tuned with just 1% of labeled data from the nuScenes-lidarseg and SemanticKITTI datasets. Our method demonstrates consistent superiority over other approaches, achieving the highest mean Intersection over Union (mIoU) scores in nearly all categories.
Additional Qualitative Results. In Figures 7, and 8, we provide additional qualitative results from the fine-tuning experiments on downstream tasks. The application of our pre-training strategies leads to a significant improvement in model performance compared to baselines initialized randomly. Particularly, our method outperforms SLiDR (Sautier et al., 2022), demonstrating its enhanced ability to handle segmentation tasks. While our approach shows substantial improvements, we observe some false positives in challenging scenarios, which we plan to address in future work.

### v2

# Junhui Hou

@@ Abstract
journal: International Journal of Computer Vision
@@ 1 Introduction
LiDAR sensors have become essential tools for capturing detailed 3D information of the environment, playing a critical role in applications such as autonomous driving, robotics, and urban planning. The rich geometric data from LiDAR point clouds provide valuable spatial awareness that is difficult to achieve with traditional 2D sensors. However, processing these 3D point clouds often requires vast amounts of labeled data to train deep neural networks effectively. Annotating point cloud data, however, is both time-consuming and expensive, posing significant challenges to the scalability and practicality of 3D deep learning models (Sautier et al., 2022; Chen et al., 2024). To address this issue, self-supervised learning has emerged as a promising solution, where networks are first trained on large volumes of unlabeled data (He et al., 2022; Caron et al., 2021). This pre-training process enables the model to learn useful feature representations without the need for costly manual annotations. Once pre-trained, the network can be fine-tuned on smaller labeled datasets, greatly reducing the need for extensive labeling efforts and improving the efficiency of training (Chen et al., 2020).
@@ 2 Related Work
3D Scene Understanding. Traditional methods for 3D scene understanding often rely on representations such as raw points (Choe et al., 2022; Chen et al., 2022a), voxels (Puy et al., 2023), range views (Kong et al., 2023a; Tian et al., 2022), and multi-view fusion (Fadadu et al., 2022; Xu et al., 2021) to capture environmental features. While effective, these approaches typically depend on large volumes of labeled data, which is both time-consuming and expensive to obtain, thereby hindering the scalability of 3D perception models (Liu et al., 2022b). To mitigate this issue, recent research has explored alternatives to reduce the reliance on fully annotated datasets. These include semi-supervised (Kong et al., 2023b; Ho et al., 2024), weakly-supervised (Liu et al., 2023; Chibane et al., 2022), self-supervised, and active learning (Luo et al., 2023; Xie et al., 2023) strategies, all of which aim to leverage minimal labeled data or automatically generate useful annotations to enhance model performance and scalability.
3D Representation Learning. Recent advancements in self-supervised learning for 3D point clouds have evolved alongside improvements in image-based methods (Zhang et al., 2024). These include pretext tasks like predicting transformations or reconstructing point cloud parts (Poursaeed et al., 2020; Sauder and Sievers, 2019). Discriminative methods focus on contrastive learning across different levels of representation (point, segment, region) to capture geometric and structural information (Nunes et al., 2022; Yin et al., 2022; Chen et al., 2022b). Temporal-consistency methods leverage spatiotemporal correlations, aligning objects over time for robust representations (Nunes et al., 2023; Huang et al., 2021). Reconstruction-based methods, such as those using Chamfer distance or surface reconstruction, aim to recover point cloud details from masked data. Cross-modal distillation approaches utilize synchronized camera-LiDAR data for contrastive learning, transferring knowledge from 2D to 3D networks (Sautier et al., 2022; Mahmoud et al., 2023; Liu et al., 2024; Chen et al., 2024; Liao et al., 2024; Puy et al., 2024; Xu et al., 2025). In this paper, we identify the limitations of these cross-distillation methods theoretically and propose a novel framework for more comprehensive 3D representation learning.
@@ 3.1 Preliminary
where $\tau$ is a temperature parameter, $M_{s}$ represents the number of sampled point-pixel pairs, and $\langle\cdot,\cdot\rangle$ is the dot product used to measure feature similarity.
@@ 4.2 Multi-modal Unified Codebook
To map the extracted modality-shared features into a unified latent space, we adopt a vector quantization (VQ) mechanism (Van Den Oord et al., 2017). This approach discretizes the continuous representations into a finite set of codewords, ensuring consistent semantic alignment across different modalities. Specifically, we define an embedding table $E=\{e_{1},e_{2},\dots,e_{V}\}\in\mathbb{R}^{V\times C}$, where $V$ is the size of the codebook, and $C$ is the dimensionality of each codeword $e_{v}$.
@@
[…]
@@ 4.5 Overall Objective Function
where reconstruction loss $\mathcal{L}_{\text{rec}}$ is the loss for the MIM task, orthogonal loss $\mathcal{L}_{\text{orth}}$ is designed to encourage independence between the modality-shared and modality-specific features within each modality, and $\mathcal{L}_{\text{kl}}$ is the Kullback-Leibler (KL) Divergence loss developed for semantic consistency proposed in OLIVINE (Zhang and Hou, 2024). The KL divergence loss is applied to the 3D features $G^{\mathrm{3D}}$.
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm, which computes the squared sum of all entries in the matrix. By minimizing the Frobenius norm of their inner product, the loss encourages these features to be uncorrelated, promoting disentangled and non-redundant representations.
@@ 5.1 Implementation Details
Pre-training Details. We pre-train the model for 50 epochs with an initial learning rate of 0.001, which is adjusted using a one-cycle learning rate scheduler (Smith, 2017) and the Adam optimizer. The 2D backbone is a Vision Transformer (ViT) pretrained with DINOv2 (Oquab et al., 2024). The pre-training tasks are performed on four RTX 3090 GPUs, with a batch size of sixteen. During pretraining, each image input is randomly masked, with 50% of its content masked at each step. For occupancy estimation, we randomly select two thousand query points per scene. For fine-tuning, we adhere to the same data splits, augmentation strategies, and evaluation protocols as those used in prior works (Mahmoud et al., 2023; Chen et al., 2024) on the nuScenes and SemanticKITTI datasets, and apply a similar procedure on other datasets.
@@ 5.2 Transfer on 3D Semantic Segmentation
Table 4: Performance of our pre-training method using the WaffleIron (WI-768) backbone. ”LP” refers to linear probing with a frozen backbone. Results are reported as mIoU on the nuScenes dataset under different proportions of labeled data.
Method LP 1% 10% 100%
No pre-training - 35.41 61.55 78.06
Ours 65.82 53.64 71.89 78.93
Furthermore, Table 4 presents the results of applying our self-supervised method to the state-of-the-art WaffleIron (WI-768) backbone (Puy et al., 2023). Under full supervision, WI-768 achieves a strong mIoU of 78.06% on the nuScenes dataset. With our self-supervised pretraining, we achieve 71.89% mIoU using only 10% of the labeled data—substantially narrowing the gap to the fully supervised performance. Notably, our method attains 65.82% mIoU through linear probing alone, which is only 12.24 points below the fully supervised result. These findings demonstrate that our approach—leveraging 2D image features without requiring manual 3D annotations—can significantly bridge the gap to fully supervised 3D representation quality.
@@ 5.5 Ablation Study
Table 7: Ablation study of each component pre-trained and fine-tuned on nuScenes. $P$: Basic pixel-point contrastive learning. Rec.: Image reconstruction and occupancy estimation tasks. Codebook: The multi-modal unified codebook. Geo.: Geometry-enhanced masked image modeling.
Exp. $P$ Rec. Codebook Geo. $\mathcal{L}_{\text{kl}}$ nuScene S.K.
@@
Figure 7: Impact of point cloud features on recovered image quality.
Effect of Geometry Enhanced Masked Image Modeling. We have included visualizations to illustrate the impact of point cloud features on the image reconstruction process. Figure 7 compares the recovered images with and without the use of point cloud features. The results show that incorporating point cloud features significantly improves the overall resemblance of the reconstructed image to the original.
Effect of 2D Backbone Settings. Thank you for your insightful comment. To evaluate the effect of masked image modeling (MIM), we conducted an ablation study with two settings: (1) freezing the 2D backbone, and (2) making the 2D backbone trainable with MIM task enabled. As shown in Table 9, allowing the 2D backbone to be updated and jointly training it with MIM leads to consistent improvements across all settings (e.g., +2.7 mIoU in linear probing, +2.2 mIoU with 5% labeled data on nuScenes). It is important to emphasize that MIM is not merely an auxiliary supervision task. As discussed in Section 3, conventional contrastive learning methods primarily focus on learning modality-shared features while neglecting modality-specific cues. The inclusion of MIM in our framework encourages the image branch to capture modality-specific features, which are complementary to the modality-shared ones. This results in richer, more comprehensive representations that improve downstream 3D task performance.
Table 9: Ablation study on the effect of 2D backbone settings, comparing frozen vs. trainable backbones and the impact of masked image modeling (MIM) on representation quality.
Frozen MIM nuScene S.K.
LP 1% 5% 1%
$\times$ 48.5 48.8 58.9 48.0
$\times$ 51.2 51.7 61.1 49.8
@@ B. Implementation Details
[…]
@@
Figure 8: Visualization of constructed modality-shared and modality-specific features, without and with the orthogonal loss $\mathcal{L}_{\text{orth}}$.
@@ C. Additional Experimental Results
Additional Quantitative Results. Tables 10 and 11 present the per-class performance of various point cloud pretraining methods, including our approach and several baseline models, all fine-tuned with just 1% of labeled data from the nuScenes-lidarseg and SemanticKITTI datasets. Our method demonstrates consistent superiority over other approaches, achieving the highest mean Intersection over Union (mIoU) scores in nearly all categories.
As shown in Table 12, incorporating image features slightly improves the 3D occupancy estimation performance (from 56.56 mIoU to 57.34 mIoU). However, while image features can provide minor improvements in the occupancy estimation task, they do not significantly contribute to 3D representation learning. Therefore, we do not leverage image features to improve the 3D occupancy estimation task in the pipeline, as our primary focus is on enhancing 3D representation learning.
Table 12: Comparison of 3D occupancy estimation performance without and with image features.
Image Features mIoU
$\times$ 56.56
57.34
Additional Qualitative Results. As shown in Figure 8, the modality-specific and modality-shared features are clearly separated when the orthogonal loss $\mathcal{L}_{\text{orth}}$ is applied. But the features are mixed without the application of the orthogonal loss, showing some overlap and correlation between the features. The results demonstrate the effectiveness of the orthogonal loss in promoting disentangled representations.
In Figures 9, and 10, we provide additional qualitative results from the fine-tuning experiments on downstream tasks. The application of our pre-training strategies leads to a significant improvement in model performance compared to baselines initialized randomly. Particularly, our method outperforms SLiDR (Sautier et al., 2022), demonstrating its enhanced ability to handle segmentation tasks. While our approach shows substantial improvements, we observe some false positives in challenging scenarios, which we plan to address in future work.